While spraying your lawn with 2-4D (an herbicide) on a windy day, about 25 g blows into your swimming pool (volume 500,000 L). The maximum acceptable concentration for 2-4D in drinking water is 0.10 ppm. Would you swim in the pool? Justify your answer
1
Expert's answer
2020-11-20T08:03:08-0500

mass of 2-4D in swimming pool = 25g

Volume of swimming pool = 500,000L


in ppm = "\dfrac{25g}{500,000L} × 1,000,000"


since 1L = 1000g,


"= \dfrac{25g}{500,000,000g} × 1,000,000=\dfrac{25}{500}"


"=0.05ppm"


Since the maximum acceptable concentration of the herbicide is 0.10ppm, it is still safe to swim in the pool, as the concentration of the swimming pool presently is 0.05ppm